If_the_sum_of_the_ages_of_a_girl_and_her_brother_is_26_years.3_years_ago_her_age_was_4_times_the_age_of_her_brotherFindout_the_ages_of_girl_and_brother
Answer by Jk22(389)

g + b = 26 y

g-3 = 4(b-3)


Subtraction of equ. 2 from equ. 1 gives an equation for b only :


b + 3 = 26 - 4b + 12 | +4b on both side

5b + 3 = 38..........| less 3

5b = 35..............| divided by 5

b = 7 y : age of brother is seven years, age of girls is 26-7 = 19 y.


Verification :

sum = 19 + 7 = 26

3 years ago : 16 = 4 * 4, 4 = 7-3.